Consider the following data:

PAE = 4,500 + 0.7Y – 25,000r

Interest rate = 6%

Y* = 11,750

 

The Fed should _____ the interest rate in order to close what type of gap.


1
Expert's answer
2021-11-30T15:00:28-0500

Planned Aggregate Expenditure (PAE)

"PAE = 4,500 + 0.7\times 11,750 – 25,000\times 0.06\\
= 11225"

PAE = 11225

The Fed should increase the interest rate in order to close the recession gap.